OpenStudy (anonymous):

Solve. 4y = 144 (Points : 1) a 36 b 16 c -16 d -36

OpenStudy (one098):

A. 36

OpenStudy (anonymous):

thanks @One098

OpenStudy (anonymous):

do u know this one Solve. -3 3 -27 27

OpenStudy (anonymous):

@One098

OpenStudy (one098):

Wait, what am I supposed solve?

OpenStudy (anonymous):

oh whoops sorry, it's -1/3b = 9

OpenStudy (anonymous):

it's negative 1 over 3b btw

OpenStudy (one098):

Its okay, and the answer is b=-9/3 or simplified, b=-3/1 or just -3.